OpenStudy (anonymous):

Solve the equation 2 over 5 x − 3 = 13 for x. Please show your work and solution in the space provided.

OpenStudy (anonymous):

plz show full work

OpenStudy (anonymous):

2 /5(40) − 3 = 13 x=40 so 2/5=0.4 0.4x40-3=13

OpenStudy (anonymous):

2/5 is dividing by doing so it makes it easier to isolate